OpenStudy (anonymous):

If you started with 20.0 grams of H2S and 50.0 grams of O2, how many grams of S8 would be produced, assuming 96% yield?

OpenStudy (anonymous):

Can you come up with the balanced chemical equation?

OpenStudy (anonymous):

Yes, one sec!

OpenStudy (anonymous):

8H2S(g)+4O2(g)=S8(l)+8H2O(g)

OpenStudy (aaronq):

you'd wanna find the limiting reactant now.

OpenStudy (anonymous):

Under optimal conditions the Claus process gives 98% yield of S8 to H2S. Does that help at all?

OpenStudy (aaronq):

you would find the limiting reactant by find the amount of moles of each reacting species and dividing it by it's stoichiometric coefficient (from the balanced equation).

OpenStudy (aaronq):

by finding*

OpenStudy (aaronq):

obviously, the one with less moles is the limiting factor.

OpenStudy (anonymous):

I'm lost.

OpenStudy (aaronq):

you would find the number of moles \(n=\dfrac{m}{M}\) M= molar mass, m=mass (in grams), n= moles then divide each by it's coefficient in the equation. 8H2S(g)+4O2(g)=S8(l)+8H2O(g) e.g. \(\dfrac{n_{O_2}}{4}\)= number of "normalized" moles i'm using the term "normalized" in a statistical sense.